# SPADE: Sparsity-Guided Debugging for Deep Neural Networks

### v1

@@ Abstract
Interpretability, broadly defined as mechanisms for understanding why and how machine learning models reach their decisions, is one of the key open goals at the intersection of deep learning theory and practice. Towards this goal, multiple tools have been proposed to aid a human examiner in reasoning about a network’s behavior in general or on a set of instances. However, the outputs of these tools—such as input saliency maps or neuron visualizations—are frequently difficult for a human to interpret, or even misleading, due, in particular, to the fact that neurons can be multifaceted, i.e., a single neuron can be associated with multiple distinct feature combinations. In this paper, we present a new general approach to address this problem, called SPADE, which, given a trained model and a target sample, uses sample-targeted pruning to provide a “trace” of the network’s execution on the sample, reducing the network to the connections that are most relevant to the specific prediction. We demonstrate that preprocessing with SPADE significantly increases both the accuracy of image saliency maps across several interpretability methods and the usefulness of neuron visualizations, aiding humans in reasoning about network behavior. Our findings show that sample-specific pruning of connections can disentangle multifaceted neurons, leading to consistently improved interpretability.
{arshia.soltanimoakhar, eugenia.iofinova, dan.alistarh}@ista.ac.at
@@ 1 Introduction
Neural network interpretability seeks mechanisms for understanding why and how deep neural networks (DNNs) make decisions, and ranges from approaches which seek to link abstract concepts to structural network components, such as specific neurons, e.g., (Erhan et al., 2009; Yosinski et al., 2015; Mordvintsev et al., ; Nguyen et al., 2016), to approaches which aim to trace individual model outputs on a per-sample basis, e.g., (Simonyan et al., 2013). While this area is developing rapidly, there is also work questioning the validity of localized explanations with respect to the model’s true decision process, pointing out confounding factors across current explainability methods and metrics (Shetty et al., 2019; Rebuffi et al., 2020; Casper et al., 2023).
@@
One key confounder for interpretability the fact the neurons of a trained, accurate DNN are often multifaceted (Nguyen et al., 2016), in the sense that they respond to many different types of features, which may be unrelated. This phenomenon directly impacts interpretability methods, such as visualizing inputs which maximize a neuron’s activation: the resulting representative input superimposes salient features, and is therefore hard to interpret. Thus, there is significant effort in the literature on addressing this issue: for instance, early work by Nguyen et al. (2016) proposed retraining the network with specialized regularizers which promote feature “disentanglement,” whereas recently Wong et al. (2021) enforced output decisions to be based on very few features by retraining the final linear output layer from scratch to be extremely sparse. Yet, one key limitation of this line of work is that generating a “debuggable” model with disentangled representations requires heavy retraining of the original model. Beyond computational cost, a conceptual issue is that the interpretations generated on top of the “debuggable” model no longer correspond to the original model’s predictions.
@@
In this paper, we propose an alternative approach called Sparsity-Guided Debugging (SPADE), which removes the above limitations, based on two main ideas: first, instead of retraining the model to become interpretable, we disentangle the feature representations for the model itself; second, this disentanglement is done for the individual sample for which we wish to obtain an interpretation. This procedure is performed efficiently, without the computational costs of retraining.
@@
Concretely, given a DNN $M$ and a sample $s$ whose output $M(s)$ we wish to interpret, SPADE functions as a pre-processing stage, in which we execute the sample $s$, together with a set of its augmentations, through the network layer-by-layer, sparsifying each layer maximally while ensuring that the output of the sparse layer still matches well with the original layer output on the sample. Thus, we obtain a sparse model $Sparse(M,s)$, which matches the original on the sample $s$, but for which extraneous connections have been removed via sample-dependent pruning. Once the custom model $Sparse(M,s)$ is obtained, we can execute any interpretability method on this subnetwork to extract a sample-specific feature visualization or saliency map. See Figure 1 for an illustration.
@@
SPADE can be implemented efficiently by leveraging solvers for accurate one-shot pruning, e.g., Frantar & Alistarh (2022), and can significantly improve performance across interpretability methods and applications. First, we illustrate SPADE by coupling it with 10 different interpretability techniques in the context of a DNN backdoor attack. Here, we find that, on a standard ResNet50/ImageNet setting, SPADE reduces the average error, taken across these methods, to less than half, from 9.91% to 4.22%. By comparison, the method of Wong et al. (2021), reduces error by 0.54% on average, in the same setup. In addition, the results of a user study we performed evaluating the impact of SPADE on the quality of feature visualization shows that, in a setting where the ground truth is determined but unknown to the user, users were significantly more successful (69.8% vs 56.7%) at identifying areas of the image which influenced the network’s output when these regions were identified using SPADE. In summary, our contributions are as follows:
Figure 1: SPADE disambiguates feature visualizations and improves the accuracy of saliency maps. This model was trained with some of the training images augmented with Trojan patches. The visualization of the ‘Albatross’ class neuron consists of a mix of natural and Trojan features, which is difficult for a human to interpret. However, preprocessing a model using a albatross image or a sample with a Trojan patch decouples the bird and fish emoji facets. Likewise, preprocessing the network with SPADE before computing a saliency map concentrates it on the Trojan patch, correctly explaining the prediction into the ‘Goose’ class. Further examples are available in Appendix G.
@@
    We provide a new interpretability-enhancing technique called SPADE, which can be applied to arbitrary models and samples to create an easier-to-interpret model “trace” customized to the specific target sample. Intuitively, SPADE works by disentangling the neurons’ superimposed feature representations via sparsification in a way that is sample-specific, which allows virtually all interpretability approaches to be more accurate.
@@
    We validate SPADE practically for image classification, by coupling it with several methods for feature visualization and generating saliency maps. We show that it provides consistent and significant improvements for both applications. Moreover, these improvements occur across all visualization methods studied, and for different models types and datasets.
@@
    We show that SPADE can be practical: it can be implemented in a computationally-efficient manner, executing in tens of minutes per instance on a single GPU. Moreover, through ablation studies, we examine the impact of task, augmentation strategy, target sample selection, and sparsity levels, showing that SPADE is robust to variations across parameters.
@@ 2 Related Work
As neural-network based models have been increasingly deployed in important or sensitive applications, there has been a corresponding increase in community and media attention to systematic errors and biases often exhibited by these systems, e.g., Buolamwini & Gebru (2018). This has led to great interest in using various techniques to aid humans in examining and debugging the models’ outputs. An overview of these approaches can be found in Linardatos et al. (2020).
One common desideratum in this space is to predict which parts of an input (e.g., image pixels) are most useful to the final prediction. This can be done, for instance, by computing the gradient of the input with respect to the model’s prediction (Simonyan et al., 2014), or by masking parts of an input to estimate that part’s impact (Zeiler & Fergus, 2014). While these techniques can be helpful in diagnosing issues, they are also prone to noisy signals (Hooker et al., 2019) and being purposefully misled (Geirhos et al., 2023). Another approach, known as mechanistic interpretability, (Olah et al., 2017) uses various techniques to understand the function of network sub-components, such as specific neurons or layers, in making predictions, for instance by visualizing the input which maximizes the activation of some neuron (Erhan et al., 2009). We emphasize that our work is not in direct competition with either of these categories of methods. Instead, our work proposes a preprocessing step to the model examination process, which should consistently improve performance.
@@
Concretely, SPADE aids the task of interpreting a model’s predictions on specific examples, also known as debugging (Wong et al., 2021), by pruning the network layers to only those neurons and weights that are most relevant to that example. Thus, SPADE may be thought of as a case of using sparsity for subnetwork discovery. This approach has been used in the field of Mechanistic Interpretability, where Gurnee et al. (2023) uses sparse linear probes to find the most relevant units to a prediction. Cao et al. (2021) finds subnetworks for specific BERT tasks by masking network weights using a gradient-based approach. Conversely, Meng et al. (2022) uses input corruption to trace out pathways in GPT models that are important for a specific example; however, their method is not based on pruning and is not evaluated in terms of interpretability metrics.
Additionally, some works aim to train sparse, and therefore more debuggable, networks. Voita et al. (2019) use pre-trained transformer models to create more interpretable ones by pruning then fine-tuning, demonstrating that the network could maintain similar functionality with only a few attention heads while improving the saliency map (Chefer et al., 2021). Other methods have focused on training more interpretable sparse models from scratch, removing the issues inherent in retraining. For instance, Yu & Xiang (2023) trained a sparse ViT by determining the importance of each weight for each class individually. Their qualitative analysis showed that their sparse model was more interpretable than dense models. Similarly, Liu et al. (2023) proposed a sparse training method inspired by the brain. This approach allowed them to identify the modular role of individual neurons in small-scale problems.
Most related, in Wong et al. (2021), the authors retrain the final fully-connected classification head of a trained network to be highly sparse, improving the attribution of predictions to the neurons in the preceding layer. This benefit arises because, after pruning, each class depends on fewer neurons from the previous layer, thus simplifying the task of individually examining connections. Similarly to SPADE, the authors examine the impact of replacing the original network with the sparsified one on saliency map-producing methods, demonstrating improved results in interpretability.
@@
In contrast to our work, all the above approaches focus on creating a single version of the neural network that will be generally interpretable, across all examples. Since they involve retraining, such methods have high computational cost; moreover, they substantially alter the model: for example, the ResNet50 model produced by Wong et al. (2021) have 72.24% ImageNet accuracy, 1.70% less than their dense baseline. Conversely, SPADE can operate on any pretrained network, and creates a customized network pruned for each target, in one-shot, which can then consistently improve performance of almost any interpretability method. Further, we demonstrate in Section 3.2 that there is a high degree of agreement between the models generated by SPADE and the original model, and in Section 4.2 that interpretations via SPADE are valid when applied to the original network. As such, SPADE is the first method which leverages sparsity to provide interpretations that are consistent with the original network.
@@ 3.1 Algorithm Overview
We now describe our method, Sparsity-Guided Debugging (SPADE). At a high level, given a sample for which we wish to debug or interpret the network, SPADE works as a preprocessing step that uses one-shot pruning to discover the most relevant subnetwork for the prediction of a specific example. We illustrate the SPADE process in Figure 2.
@@
We start with an arbitrary input sample chosen by the user, which we would like to interpret. SPADE then expands this sample to a batch of samples by applying augmentation techniques. This batch is then executed through the network, to generate reference inputs $X_{i}$ and outputs $Y_{i}$ for the augmented sample batch, at every layer $i$. Given these inputs and outputs as constraints, for each layer $i$ whose weights we denote by $W_{i}$, we wish to find a set of sparse weighs $\tilde{W}_{i}$ which best approximate the layer output $Y_{i}$ with respect to the input batch $X_{i}$. In our implementation, we adopt the $\ell_{2}$ distance metric. Thus, for a linear layer, we would like to find
$\tilde{W}_{i}=\textnormal{argmin}_{W\textnormal{sparse}}\|WX_{i}-Y_{i}\|_{2}^{2}.$ (1)
@@
To solve this constrained optimization problem at each layer, we use a custom sparsity solver (Frantar & Alistarh, 2022). We discuss specific implementation details in the next section.
@@
Once layer-wise pruning has completed, we have obtained a model that has been pruned specifically relative to our target sample and its augmentations. Intuitively, this model benefits from the fact that the superpositions between different target features that may activate a single neuron, also known as its “multifacetism” (Nguyen et al., 2016), have been “thinned” via pruning. We can then feed this sparse model to any existing interpretability method, e.g., Sundararajan et al. (2017); Zeiler & Fergus (2014); Olah et al. (2017). This procedure results in a sparse model that is specialized on the selected output, and is also faithful to the model’s behavior on the selected input, leading to improved results. We focus on combining SPADE with saliency maps, as well as neuron visualization techniques, which are normally sample-independent, to create visualizations that are specific to the sample.
1:procedure SPADE Algorithm($M,s,I$)
2:     $\triangleright$ $M$: Model, $s$: Sample, $I$: Interpretability Method
3:     $B\leftarrow$ Empty $\triangleright$ Batch of Augmented samples
4:     for  Augmentation Batch Size  do
5:         Append a random augmentation of $s$ to $B$
6:     end for
7:     for  Each layer in $M$  do

### v2

@@ Abstract
It is known that sparsity can improve interpretability for deep neural networks. However, existing methods in the area either require networks that are pre-trained with sparsity constraints, or impose sparsity after the fact, altering the network’s general behavior. In this paper, we demonstrate, for the first time, that sparsity can instead be incorporated into the interpretation process itself, as a sample-specific preprocessing step. Unlike previous work, this approach, which we call SPADE, does not place constraints on the trained model and does not affect its behavior during inference on the sample. Given a trained model and a target sample, SPADE uses sample-targeted pruning to provide a “trace” of the network’s execution on the sample, reducing the network to the most important connections prior to computing an interpretation. We demonstrate that preprocessing with SPADE significantly increases the accuracy of image saliency maps across several interpretability methods. Additionally, SPADE improves the usefulness of neuron visualizations, aiding humans in reasoning about network behavior. Our code is available at https://github.com/IST-DASLab/SPADE.
Machine Learning, ICML
[…]
@@ 1 Introduction
Neural network interpretability seeks mechanisms for understanding why and how deep neural networks (DNNs) make decisions, and ranges from approaches which seek to link abstract concepts to structural network components, such as specific neurons, e.g., (Erhan et al., 2009; Yosinski et al., ; Mordvintsev et al., ; Nguyen et al., 2016), to approaches which aim to trace individual model outputs on a per-sample basis, e.g., (Simonyan et al., 2013). While this area is seeing a lot of interest, there is also work questioning the validity of localized explanations with respect to the model’s true decision process, pointing out confounders across current explainability methods and metrics (Shetty et al., 2019; Rebuffi et al., 2020; Casper et al., 2023).
Figure 1: Given an input image and model, SPADE prunes the model using image augmentations. The resulting trace (subnetwork) can be used with existing interpretability methods to increase their usefulness and accuracy.
Figure 2: SPADE disambiguates feature visualizations and improves the faithfulness of saliency maps. (Left) The ”Lighter, igniter” class neuron visualization does not give useful clues for why the Matchstick and Spotlight images were incorrectly classified into that class. The visualizations obtained with SPADE identify a matchstick head pattern in the first case and a flame pattern in the second case, suggesting that these may be spurious features for the Lighter class. (Right) A model implanted with Trojan patches leads to a Fox image being misclassified as a Goose. In this case, we are confident that the heart emoji was entirely responsible for the misclassification - yet, the saliency map without SPADE incorrectly assigns large saliency scores to large parts of the fox image. Conversely, the saliency map obtained with SPADE correctly identifies the emoji pixels. Best viewed in color. Further examples are available in Appendix J.
@@
One key confounder for interpretability is the fact the neurons of a trained, accurate DNN often respond to many different types of features, which may be unrelated (Nguyen et al., 2016; Olah et al., 2020, 2017). For example,  Olah et al. (2017) finds a neuron equally likely to respond to car shields and cat paws, and with the same intensity. This phenomenon directly impacts interpretability methods, such as visualizations of inputs that maximize a neuron’s activation: the resulting representative input superimposes salient features, and is therefore hard to interpret. Thus, there is significant effort in the literature on addressing this issue: for instance, early work by Nguyen et al. (2016) proposed retraining the network with specialized regularizers which promote feature “disentanglement,” whereas Wong et al. (2021) enforced output decisions to be based on very few features by retraining the final linear output layer from scratch to be extremely sparse. Yet, one key limitation of this line of work is that generating a “debuggable” model with disentangled representations requires heavy retraining of the original model, which may be impractical or impossible. Beyond cost, a conceptual issue is that the interpretations generated on top of the retrained “debuggable” model no longer correspond to the original model’s predictions.
@@
We propose an alternative approach called Sparsity-Guided Debugging (SPADE), which removes the above limitations, based on two main ideas: first, instead of retraining the model to become interpretable, we disentangle the feature representations for the original model; second, this disentanglement is done for the individual sample for which we wish to obtain an interpretation. This procedure can be performed efficiently, without the computational costs of retraining.
@@
We illustrate the process in Figure 1. Given a DNN $M$ and a sample $s$ whose output $M(s)$ we wish to interpret, SPADE functions as a pre-processing step, in which we execute the sample $s$, together with a set of its augmentations, through the network layer-by-layer, sparsifying each layer while ensuring that the output of the sparse layer still matches well with the original layer output on the sample. Thus, we obtain a sparse trace $Sparse(M,s)$, which matches the original on the sample $s$, but for which extraneous connections relative to this sample’s output have been removed via sample-dependent pruning. Once the custom trace $Sparse(M,s)$ is obtained, we can execute any interpretability method on this pruned network to extract a sample-specific feature visualization or saliency map.
@@
We show that SPADE can be implemented efficiently by leveraging solvers for accurate one-shot pruning (Frantar & Alistarh, 2022, 2023), and can significantly improve performance across interpretability methods and applications (Figure  2). First, we illustrate SPADE by coupling it with 10 different saliency map creation techniques. In the context of a DNN backdoor attack (Figure 2, right panel), we find that, in a standard ResNet50/ImageNet setting, SPADE reduces the average error, taken across all methods, to less than half, from 8.99% to 3.45%. By comparison, the prior method of (Wong et al., 2021), reduces error by 0.49% on average, in the same setup. Additionally, we demonstrate that SPADE increases the fidelity of input attribution methods by measuring the impact of SPADE on standard insertion and deletion metrics, where we model confidence is measured when the most salient input components (e.g., pixels) are added or removed, respectively. This test further validates our claim that interpretations formed with the aid of SPADE apply to the original, dense model.
Further, the results of a human user study we performed, evaluating the impact of SPADE on the quality of feature visualization, shows that, in a setting where the ground truth is determined but unknown to the user, users were significantly more successful (69.8% vs 56.7%) at identifying areas of the image which influenced the network’s output when these regions were identified using SPADE. In summary, our contributions are as follows:
@@
    We demonstrate, for the first time, that post-hoc sample-specific sparsification aids interpretability for pretrained models, without requiring sparsity to be imposed during the training or inference process.
@@
    We provide a new interpretability-enhancing technique called SPADE, which can be applied to arbitrary models and samples to create an easier-to-interpret model “trace” customized to the specific target sample. Intuitively, SPADE disentangles the neurons’ superimposed feature representations in a way that is sample-specific, which allows virtually all interpretability approaches to be more accurate with respect to the dense model.
@@
    We validate SPADE practically for image classification, by coupling it with methods for feature visualization and saliency map generation. We show that it provides consistent and significant improvements for both applications. Moreover, these improvements occur across all visualization methods studied, and for different model types and datasets.
  4. 4.
    We show that SPADE can be practically implemented in a computationally-efficient manner. In its fastest version, SPADE requires approximately 3 seconds per sample for a ResNet50 model on a single GPU, enabling it to be run interactively. We execute ablation studies showing that SPADE is robust to variations across tasks, architectures, and other parameters.
@@ 2 Related Work
As DNN-based models are increasingly deployed in important or sensitive applications, there has been an increase in attention to systematic errors and biases often exhibited by these systems, e.g., Buolamwini & Gebru (2018). This has led to interest in aiding humans in examining and debugging the models’ outputs. An overview of the area can be found in (Linardatos et al., 2020).
One common desideratum in this space is to predict which parts of an input (e.g., image pixels) are most useful to the final prediction. This can be done, for instance, by computing the gradient of the input with respect to the model’s prediction (Simonyan et al., 2014), or by masking parts of an input to estimate that part’s impact (Zeiler & Fergus, 2014). While these techniques can be helpful in diagnosing issues, they are also prone to noisy signals (Hooker et al., 2019) and being purposefully misled (Geirhos et al., 2023), and, in the case of linear methods, have provable limits on generalization (Bilodeau et al., 2022). Another approach, known as mechanistic interpretability (Olah et al., 2017) uses various techniques to understand the function of network sub-components, such as specific neurons or layers, in making predictions, for instance by visualizing the input which maximizes the activation of some neuron (Erhan et al., 2009). We emphasize that our work is not in direct competition with either of these categories of methods. Instead, our work proposes a preprocessing step to the model examination, which consistently improves performance.
@@
Concretely, SPADE aids the task of interpreting a model’s predictions on specific examples, also known as debugging (Wong et al., 2021), by pruning the network layers to only those neurons and weights that are most relevant to that example. Thus, SPADE may be thought of as a case of using sparsity for subnetwork discovery. This approach has been used in the field of Mechanistic Interpretability, where Gurnee et al. (2023) used sparse linear probes to find the most relevant units to a prediction. Cao et al. (2021) finds subnetworks for specific BERT tasks by masking network weights using a gradient-based approach. Conversely, Meng et al. (2022) uses input corruption to trace out pathways in GPT models that are important for a specific example and (O’Mahony et al., 2023) uses input clustering to disentangle neuron representations; however, these methods are not based on sparsity and are not evaluated in terms of interpretability metrics.
More recently, works such as linear probing  (Belrose et al., 2023; Pal et al., 2023; Wang et al., 2023) and activation patching  (Geiger et al., 2020; Kramar et al., 2024) aimed at discovering feature representation in transformer models. However, these approaches are orthogonal to existing methods such as saliency maps and cannot be combined with them. Works such as  Huben et al. (2024); Scherlis et al. (2022) take steps toward resolving polysemanticity in neurons by means of discovering individual features, a very promising line of work that may come to be complimentary to the work we present here.
Sparsity for interpretability.
Some works aim to train sparse, and therefore more debuggable, networks. Voita et al. (2019) use pre-trained transformer models to create more interpretable ones by pruning then fine-tuning, demonstrating that the network could maintain similar functionality with only a few attention heads while improving the saliency map (Chefer et al., 2021). Other methods have focused on training more interpretable sparse models from scratch, removing the issues inherent in retraining. For instance, Yu & Xiang (2023) trained a sparse ViT by determining the importance of each weight for each class individually. Their qualitative analysis showed that their sparse model was more interpretable than dense models. Liu et al. (2023) proposed a sparse training method inspired by the brain, which allowed them to identify the role of individual neurons in small-scale problems. Finally,  Panousis et al. (2023) trained interpretable sparse linear concept discovery models.
Most related, Wong et al. (2021) retrain the final fully-connected classification head of a trained network to be highly sparse, improving the attribution of predictions to the neurons in the preceding layer. This benefit arises because, after pruning, each class depends on fewer neurons from the previous layer, thus simplifying the task of individually examining connections. Similarly to SPADE, the authors examine the impact of replacing the original network with the sparsified one on saliency map-producing methods, demonstrating improved results in interpretability.
@@
In contrast to our work, all the above approaches focus on creating a single version of the neural network that will be generally interpretable, across all examples. Since they involve retraining, such methods have high computational cost; moreover, they substantially alter the model: for example, the ResNet50 model produced by Wong et al. (2021) have 72.24% ImageNet accuracy, 1.70% less than their dense baseline. We show, for the first time, that example-specific pruning can aid model interpretability and propose a method that can operate on any pretrained network, and consistently improves performance across interpretability methods. We demonstrate in Sections  4.1 and 4.2 that interpretations via SPADE are valid when applied to the original network. As such, SPADE is the first method that leverages sparsity to provide interpretations that are consistent with the original network.
@@ 3.1 Algorithm Overview
At a high level, given a sample for which we wish to debug or interpret the network, SPADE works as a preprocessing step that uses one-shot pruning to discover the most relevant subnetwork for the prediction of a specific example. We illustrate the SPADE process in Figure 1 and provide the exact algorithm in Algorithm 1.
@@
We start with an arbitrary input sample chosen by the user, which we would like to interpret. SPADE then expands this sample to a batch of samples by applying augmentation techniques [^1] Augmenting the samples in this way can influence the top-1 prediction. However, this does not affect the method, as it is prediction-agnostic.. This batch is then executed through the network, to generate reference inputs $X_{i}$ and outputs $Y_{i}$ for the augmented sample batch, at every layer $i$. Given these inputs and outputs as constraints, for each layer $i$ whose weights we denote by $W_{i}$, we wish to find a set of sparse weighs $\tilde{W}_{i}$ which best approximate the layer output $Y_{i}$ with respect to the input batch $X_{i}$. In our implementation, we adopt the $\ell_{2}$ distance metric. Thus, for a linear layer of size K and sparsity target S, we seek to find to find
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{W}_{i}=\textnormal{argmin}_{W:\|W\|_{0}\leq K\cdot S}\|WX_{i}-Y_{i}\|_{2}^{2}.}$ (1)
@@
To solve this constrained optimization problem at each layer, we use custom sparsity solvers. We discuss implementation details in the next section.
@@
Once layer-wise pruning has completed, we have obtained a trace of the target sample through the network. Intuitively, this trace benefits from the fact that the superpositions between different target features that may activate a single neuron, also known as its “polysemanticism” (Olah et al., 2020), have been “thinned” via pruning, and we therefore retain the features that are relevant to the specific input. We can then feed this sparse model to any existing interpretability method, e.g., (Sundararajan et al., 2017a; Zeiler & Fergus, 2014; Olah et al., 2017). This procedure results in a sparse model that is specialized for and faithful to the model’s behavior on the selected input. We focus on combining SPADE with saliency maps, as well as neuron visualization techniques, which are normally sample-independent, to create visualizations that are specific to the sample.
3.2 Implementation Details
Pruning approach.
The pruning approach must be chosen with care, as pruning can significantly alter the network circuitry and the predictions (Peste et al., 2021). We require that the pruning be done in a way that preserves the model’s output (by requiring that sparse outputs closely match the dense ones for each layer), and be done one-shot, without retraining. For this, one can use one of the existing one-shot sparsity solvers, e.g. (Hubara et al., 2021; Frantar & Alistarh, 2023, 2022; Kuznedelev et al., 2023). We focus on two solvers. The OBC solver (Frantar & Alistarh, 2022), provides the best approximate solution to the constrained problem in Equation 1; however, it is compute-intensive. To mitigate this, we also examine the faster but less precise SparseGPT solver (Frantar & Alistarh, 2023), which can perform the pruning procedure in about 23 seconds/sample, at the cost of low accuracy loss. This is practical for large-scale use, as we demonstrate by running the evaluation on 21 121 images in Appendix E.
  Procedure SPADE Algorithm($M,s,I$)
   {$M$: Model, $s$: Sample, $I$: Interpretability Method}
  $B\leftarrow$ Empty {Batch of Augmented samples}
  for  Augmentation Batch Size  do
      Append a random augmentation of $s$ to $B$
  end for
  for  Each layer in $M$  do